Q1. Archie runs a small mineral exploration business (as a sole proprietorship). In 2017, he purchased land (for $84,000) where he suspected a magnesium deposit was located. He incurred $18,000 of exploration costs related to the development of the magnesium mine in 2017 and an additional $32,000 of exploration costs in 2018. Archie elected to deduct these expenditures for regular tax purposes in both 2017 and 2018.

In 2019, given a decline in magnesium prices, Archie chose to sell the property for $72,000, rather than bring the mine to production.

  • AMT adjustment is required for mine exploration expenditures?
  • AMT adjustment for the sale of the land?

Part B. In 2019, Liza exercised an incentive stock option that had been granted to her in 2017 by her employer, Weather Corporation. Liza acquired 100 shares of Weather stock for the option price of $7,800 per share. The rights in the stock become freely transferable and not subject to a substantial risk of forfeiture in 2019. The fair market value of the stock at the date of exercise was $9,500 per share. Liza sells the stock for $10,700 per share later in 2021.

If an amount is zero, enter "0".

a. What is the amount of Liza's AMT adjustment in 2019? What is her recognized gain on the sale for regular income tax and for AMT purposes in 2021?

  • AMT adjustment of $?
  • Her recognized gain on the sale for regular income tax is? and  for AMT purposes?
  • As a result, Liza has a AMT adjustment in 2021?

b. How would your answers in (a) change if Liza had sold the stock in 2019 rather than 2021?

  • Liza has a AMT adjustment is required in 2019?
  • For regular tax purposes, Liza would recognize of ordinary/compensation income? and of short-term capital gains?
Answer #1

ANSWER:- (A) Archie Run A small business

Land value of purchase price-$84000

Exploration Cost(in 2017) =$18000

Exploration cost (in 2018)=$32000

Total sales valus are =$72000

-Land purchase value are Divide in Two year ,that is $84000/2=$42000

-per year cost are=$42000+$18000=$60000

-that is second year cost are=$42000+$32000=$74000

-Total sale prise-$72000

Total cost are-$74000

=profit and Loss are-2000 $ Loss

=The management suggession give the realy You think better second option

frist option in very costelly

(B)My answer will be chnange in share sales in 2021

toatl shares value in 2019=$9500

No of shares=100

net value in shares =$950000

2021 shares value in =$10700

No of shares =100

net value in share=$1070000

=value increase in one Year=1070000-950000=$120000

The best answer in share value are in 2021,

and better value are recived in shares  

i will suggess to management share hold in one year more then value in $120000 recived,